# Synthesizing Efficient and Permissive Programmatic Runtime Shields for Neural Policies

### v1

@@ Abstract.
With the increasing use of neural policies in control systems, ensuring their safety and reliability has become a critical software engineering task. One prevalent approach to ensuring the safety of neural policies is to deploy programmatic runtime shields alongside them to correct their unsafe commands. However, the programmatic runtime shields synthesized by existing methods are either computationally expensive or insufficiently permissive, resulting in high overhead and unnecessary interventions on the system. To address these challenges, we propose Aegis, a novel framework that synthesizes lightweight and permissive programmatic runtime shields for neural policies. Aegis achieves this by formulating the seeking of a runtime shield as a sketch-based program synthesis problem and proposing a novel method that leverages counterexample-guided inductive synthesis and Bayesian optimization to solve it. To evaluate Aegis and its synthesized shields, we use four representative control systems and compare Aegis with the current state-of-the-art. Our results show that the programmatic runtime shields synthesized by Aegis can correct all unsafe commands from neural policies, ensuring that the systems do not violate any desired safety properties at all times. Compared to the current state-of-the-art, Aegis’s shields exhibit a 2.1$\times$ reduction in time overhead and a 4.4$\times$ reduction in memory usage, suggesting that they are much more lightweight. Moreover, Aegis’s shields incur an average of 1.6$\times$ fewer interventions than other shields, showing better permissiveness.
@@
copyright: acmlicensedjournal: TOSEMccs: Software and its engineering Formal software verificationccs: Computing methodologies Computational control theory
@@ 1. Introduction
Recent years have seen a notable increase in the use of deep neural networks (DNNs) to support control systems like autonomous vehicles (Song et al., 2022; Ahmed et al., 2021; Lopez et al., 2023). In these systems, DNNs, also referred to as neural policies (Zhu et al., 2019; Donti et al., 2021), process information from the physical environment and send commands to hardware components like actuators to regulate their behavior in accordance with desired specifications. Neural policies excel at handling complex control tasks, such as autonomous driving (Kiran et al., 2021; Muhammad et al., 2020), aerospace guidance (Chai et al., 2021; Dong et al., 2021), and even nuclear fusion (Degrave et al., 2022; Seo et al., 2024), often outperforming traditional control software (Song et al., 2022; Lillicrap et al., 2016; Duan et al., 2016; Rajeswaran et al., 2017). According to Microsoft’s projection, by 2025, 50% of industrial companies will embrace control systems with neural policies, increasing productivity by 20% and generating \$3.7 trillion in revenue (Microsoft, 2020).
While neural policies have proven effective in various domains, concerns about their safety have arisen: they can occasionally produce unexpected or inappropriate commands, causing system failures that violate desired safety specifications (Pang et al., 2022; Amodei et al., 2016; He et al., 2024; Pei et al., 2017; Song et al., 2022; Zhang et al., 2020). This risk is exemplified in safety-critical tasks, e.g. autonomous driving, where neural policies may fail to control vehicles at desired heading angles, resulting in scenarios such as vehicles veering off the road or even colliding with pedestrians. Such catastrophic accidents have been reported multiple times in real-world systems employing neural policies, such as Tesla Autopilot (Post, 2023; Journalism, 2023) and Waymo Driver (Verge, 2023; TechCrunch, 2023). In addition, certain neural policies in aerospace and manufacturing control systems have also demonstrated unsafe behavior (SEA, 2019; Telegraph, 2023), further highlighting the safety risks associated with using neural policies in safety-critical domains.
@@
Programmatic Runtime Shields.   To date, programmatic runtime shields have demonstrated effectiveness in ensuring the safety of neural policies (Zhu et al., 2019; Alshiekh et al., 2018; Xiong and Jagannathan, 2021; Guo et al., 2023). These shields are typically simple and verifiable programs that can run alongside the neural policy and promptly correct any inappropriate commands from the policy as it executes. Figure 1 illustrates how a programmatic runtime shield works within a control system using a neural policy. Specifically, the shield monitors commands from the neural policy and intervenes when it detects potentially unsafe commands that could cause the system to violate required safety properties (line 2, marked in green). Upon intervention, the shield corrects unsafe commands by replacing them with safe alternatives based on the current state of the physical environment (lines 3-4, marked in red). This ensures that the desired safety properties are maintained at all times.
@@
Given that the synthesis of these shields is typically less costly than formal verification or exhaustive testing of neural policies, they prove to be practical in ensuring the safety of neural policies (Zhu et al., 2019; Könighofer et al., 2020; Xiong and Jagannathan, 2021). Additionally, the synthesis and operation of runtime shields usually require only the inputs and outputs of the policies, rather than the neural network parameters. This makes deploying runtime shields often the only viable option for correcting safety violations of neural policies, especially when retraining neural networks is prohibitively expensive or when accessing the parameters of the neural policies is restricted in the case of using proprietary products from third-party vendors.
@@
What Makes Good Programmatic Runtime Shields?  Three critical attributes define the efficacy of programmatic runtime shields: effectiveness, efficiency, and permissiveness (Zhu et al., 2019; Bloem et al., 2015; Könighofer et al., 2017; Wu et al., 2019; Könighofer et al., 2020). First, a good shield should effectively correct any unsafe command from the neural policy, thereby ensuring the perpetuation of desired safety properties in the system. Second, as an additional component running alongside the neural policy, the shield should be efficient, minimizing time and memory overhead. Lastly, maintaining permissiveness is essential (Zhu et al., 2019; Könighofer et al., 2023, 2020). The shield’s judgment of whether a command from the neural policy is unsafe is not always accurate, and it may intervene unnecessarily when it encounters safe commands. The shield should be permissive, minimizing unnecessary intervention and allowing the neural policy to operate as freely as possible.
@@
Deficiencies in any of these key attributes can have adverse consequences. For example, an inefficient runtime shield can consume excessive system resources, resulting in degraded system performance or crashes. This is particularly problematic in real-time systems (Kopetz and Steiner, 2022) like autonomous vehicles, where timely responses are critical, or in resource-constrained embedded systems (Roth et al., 2024; Panda et al., 1999) like drones, where memory is limited. In addition, a non-permissive shield can often interfere with safe commands, thereby degrading system stability (i.e., a system’s ability to achieve steady states in response to random inputs or disturbances, typically measured by required time steps (Zhu et al., 2019); see Section 5.1 for further discussion). Unfortunately, current state-of-the-art programmatic runtime shields still have limitations with respect to these key attributes, as we discuss next.
@@
Several methods for synthesizing programmatic runtime shields have emerged in recent years, among which Zhu et al.’s work (An Inductive Synthesis Framework for Verifiable Reinforcement Learning (Zhu et al., 2019), hereafter referred to as VRL) stands out as the state-of-the-art. VRL crafts programmatic runtime shields capable of effectively rectifying all unsafe commands to consistently uphold desired safety properties (see Section 2 for more details). However, these shields have limitations in the other two attributes. First, VRL’s shields involve operations with multiple high-degree polynomials, resulting in a relatively large time and memory overhead on the systems they operate within. Second, their lack of permissiveness often leads to unnecessary interventions in safe commands, consequently compromising system stability. These limitations highlight opportunities for improving VRL’s shields, necessitating our work to extend the state-of-the-art.
@@
We introduce Aegis, a novel method to synthesize efficient and permissive programmatic runtime shields for neural policies. The core idea of Aegis is to formulate the seeking of a shield as a sketch-based program synthesis problem and use counterexample-guided inductive synthesis (Solar-Lezama, 2008) and Bayesian optimization (Garnett, 2023) to solve it. These two techniques have been studied in program synthesis in the past (Abate et al., 2018; Češka et al., 2019; Saad et al., 2019). However, Aegis is the first to combine them to enhance the efficiency and permissiveness of runtime shields, offering a unique perspective to the field.
@@
Two thrusts make Aegis superior to the state-of-the-art. First, the shields synthesized by Aegis involve only linear functions instead of the high-degree polynomials in VRL’s shields. This design choice reduces computational complexity and overhead, making our shields much more efficient. Second, our Bayesian optimization algorithm assigns the optimal threshold to the switching strategy by minimizing unnecessary interventions, thus making our shields more permissive than those synthesized by VRL.
@@
Our results show that the programmatic runtime shields synthesized by Aegis effectively correct all unsafe commands from neural policies, ensuring that the systems do not violate the desired safety properties at any time. Notably, compared to VRL, Aegis’s shields exhibit a 2.1$\times$ reduction in time overhead and a 4.4$\times$ reduction in memory overhead, suggesting they are much more efficient. Aegis’s shields also show improved permissiveness, incurring an average of 1.6$\times$ fewer interventions than those synthesized by VRL. We also gauge the ratio of necessary interventions to all interventions and find that, on average, our shields intervene when necessary 9.9% more frequently than VRL’s. In addition, Aegis shows superior scalability relative to VRL, reducing synthesis time by an average of 77.3%. Additionally, we provide an analysis of replacing neural policies with Aegis’s shields, confirming the necessity for the coexistence of neural policies and shields to maintain the system’s safe and stable operation.
Main Contributions
Our main contributions are:
  • We improve the efficiency as well as permissiveness of programmatic runtime shields for neural policies, addressing the limitations of the state-of-the-art.
  • We propose and implement Aegis, a novel method that synthesizes programmatic runtime shields via a sketch-based program synthesis approach.
  • We evaluate Aegis on four representative systems and show its superiority over the state-of-the-art, with significant improvements in all metrics.
@@ 2. Preliminary
Modern control systems are typically implemented with digital rather than analog techniques (Lee, 2015; Zhu et al., 2019), where signals are processed discretely over fixed time intervals. Thus, the state transition of an environment can be formulated with discrete and fixed time step $t$ using Euler’s method (Hildebrand, 1987; Zhu et al., 2019). Namely, a state $s_{t}$ transitions to the subsequent state $s_{t+1}$ after time $t$ by the following equation:
@@
Here, $c_{t}$ denotes a command generated by a neural policy $\pi$, which is implemented as a neural network that takes an environment state $s_{t}$ as input, i.e., $c_{t}=\pi(s_{t})$, and $\Delta t$ is the fixed time step. Typically, neural policies are trained using Deep Reinforcement Learning (DRL) algorithms (Li, 2017; Lillicrap et al., 2016), which optimize the policy’s parameters to maximize the expected discounted cumulative reward through interaction with the environment. Specifically, in an environment $\mathcal{E}$, a system begins at an initial state $s_{0}\in\mathcal{S}_{0}$, and its neural policy produces commands to control the system to interact with the environment until it reaches a termination state, $s_{T}$, after $T$ time steps. The interaction history is called a trajectory:
@@
The specific DRL algorithms for solving the above equation vary with different tasks and are beyond our paper’s scope. We refer interested readers to (Li, 2017; Lillicrap et al., 2016) for more details.
1def main_loop():
2 $\eta$, $\omega$ = read_environment()
[…]
5 motor.update($c$)
Figure 2. Main control loop in a quadcopter environment (simplified for readability). The green and red colors highlight the neural policy and runtime shield, respectively.
@@ 2 $\eta$, $\omega$ = read_environment()
1def programmatic_runtime_shield($c$, $\eta$, $\omega$):
2 $\hat{\eta}$, $\hat{\omega}$ = environment_model($c$, $\eta$, $\omega$)
[…]
5 else:
6 return $c$
@@ 2 $\hat{\eta}$, $\hat{\omega}$ = environment_model($c$, $\eta$, $\omega$)
1def programmatic_runtime_shield($c$, $\eta$, $\omega$):
[…]
4 return $\kappa$
5 else:
6 return $c$
Figure 4. A runtime shield synthesized by Aegis for the neural policy in a quadcopter system. Line 2 (marked in red) is a linear policy that outputs safe commands, while line 3 (marked in green) identifies potentially unsafe commands. Line 4 returns the safe command if the one from the neural policy is potentially unsafe.
@@ 3. Aegis
This section provides details of Aegis, with its workflow illustrated in Figure 5. Aegis begins by constructing a program sketch that consists of a linear policy and a switching strategy (Section 3.1). Then, given a complex and possibly unknown physical environment, Aegis derives its state transition function by numerical differentiation (Griewank, 2013) (Section 3.2). Next, Aegis initializes a linear policy using the Linear-Quadratic Regulator (LQR) algorithm (Anderson and Moore, 2007) and iteratively refines it with a counterexample-guided inductive synthesis algorithm, until the linear policy is verified to be safe against the safety specifications (Section 3.3), determining the parameters of the linear policy. Following this, Aegis synthesizes the threshold of the switching strategy using a Bayesian optimization algorithm, aiming to minimize unnecessary shield interventions (Section 3.4). Finally, the synthesized components are incorporated into the program sketch to form a programmatic runtime shield. The following subsections describe each step in detail.

### v3

@@ Abstract.
[…]
@@
copyright: ccjournal: TOSEMjournalyear: 2025journalvolume: 1journalnumber: 1article: 1publicationmonth: 1doi: 10.1145/3773034ccs: Software and its engineering Software verificationccs: Software and its engineering Automatic programmingccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Recent years have seen a notable increase in the use of deep neural networks (DNNs) to support control systems like autonomous vehicles (Song et al., 2022; Ahmed et al., 2021; Lopez et al., 2023). In these systems, DNNs, also referred to as neural policies (Zhu et al., 2019; Donti et al., 2021), process information from the physical environment and send commands to hardware components like actuators to regulate their behavior in accordance with desired specifications. Neural policies excel at handling complex control tasks, such as autonomous driving (Kiran et al., 2022; Muhammad et al., 2021), aerospace guidance (Chai et al., 2021; Dong et al., 2021), and even nuclear fusion (Degrave et al., 2022; Seo et al., 2024), often outperforming traditional control software (Song et al., 2022; Lillicrap et al., 2016; Duan et al., 2016; Rajeswaran et al., 2017). According to Microsoft’s projection, by 2025, 50% of industrial companies will embrace control systems with neural policies, increasing productivity by 20% and generating \$3.7 trillion in revenue (Microsoft, 2020).
While neural policies have proven effective in various domains, concerns about their safety have arisen: they can occasionally produce unexpected or inappropriate commands, causing system failures that violate desired safety specifications (Pang et al., 2022; Amodei et al., 2016; Raji and Dobbe, 2023; He et al., 2024; Pei et al., 2017; Song et al., 2022; Zhang et al., 2020). This risk is exemplified in safety-critical tasks, e.g., autonomous driving, where neural policies may fail to control vehicles at desired heading angles, resulting in scenarios such as vehicles veering off the road or even colliding with pedestrians. Such catastrophic accidents have been reported multiple times in real-world systems employing neural policies, such as Tesla Autopilot (Post, 2023; Journalism, 2023) and Waymo Driver (Verge, 2023; TechCrunch, 2023). In addition, certain neural policies in aerospace and manufacturing control systems have also demonstrated unsafe behavior (SEA, 2019; Telegraph, 2023), further highlighting the safety risks associated with using neural policies in safety-critical domains.
@@
To date, programmatic runtime shields have demonstrated effectiveness in ensuring the safety of neural policies (Zhu et al., 2019; Alshiekh et al., 2018; Xiong and Jagannathan, 2022; Guo et al., 2023). These shields are typically simple and verifiable programs that can run alongside the neural policy and promptly correct any inappropriate commands from the policy as it executes. Figure 1 illustrates how a programmatic runtime shield works within a control system using a neural policy. Specifically, the shield monitors commands from the neural policy and intervenes when it detects potentially unsafe commands that could cause the system to violate required safety properties (line 2, marked in green). Upon intervention, the shield corrects unsafe commands by replacing them with safe alternatives based on the current state of the physical environment (lines 3-4, marked in red). This ensures that the desired safety properties are maintained at all times.
@@
Given that synthesizing these shields is generally less expensive than performing formal verification or exhaustive testing of neural policies, they offer a practical approach to ensuring their safety (Zhu et al., 2019; Könighofer et al., 2020; Xiong and Jagannathan, 2022). Moreover, both the synthesis and operation of runtime shields typically rely only on the inputs and outputs of the policies, without requiring access to the neural network parameters. As a result, deploying runtime shields is often the only viable option for addressing safety violations of neural policies, particularly when retraining neural networks is prohibitively expensive or when accessing the parameters of the neural policies is restricted in the case of using proprietary products from third-party vendors.
@@
What Makes Good Programmatic Runtime Shields?  Three critical attributes define the efficacy of programmatic runtime shields: effectiveness, efficiency, and permissiveness (Zhu et al., 2019; Bloem et al., 2015; Könighofer et al., 2017; Wu et al., 2019; Könighofer et al., 2020). First, a good shield should effectively correct any unsafe command from the neural policy, thereby ensuring the perpetuation of desired safety properties in the system. Second, as an additional component running alongside the neural policy, the shield should be efficient, minimizing time and memory overhead. Lastly, maintaining permissiveness is essential (Zhu et al., 2019; Könighofer et al., 2022, 2020). The shield’s judgment of whether a command from the neural policy is unsafe is not always accurate, and it may intervene unnecessarily when it encounters safe commands. The shield should be permissive, minimizing unnecessary intervention and allowing the neural policy to operate as freely as possible.
@@
Deficiencies in any of these key attributes can have adverse consequences. For example, an inefficient runtime shield can consume excessive system resources, resulting in degraded system performance or crashes. This is particularly problematic in real-time systems (Kopetz, 2011) like autonomous vehicles, where timely responses are critical, or in resource-constrained embedded systems (Roth et al., 2024; Panda et al., 1998) like drones, where memory is limited. In addition, a non-permissive shield can often interfere with safe commands, thereby degrading system stability (i.e., a system’s ability to achieve steady states in response to random inputs or disturbances, typically measured by required time steps (Zhu et al., 2019); see Section 5.1 for further discussion). Unfortunately, current state-of-the-art programmatic runtime shields still have limitations with respect to these key attributes, as we discuss next.
@@
Several methods for synthesizing programmatic runtime shields have emerged in recent years, among which Zhu et al.’s work (An Inductive Synthesis Framework for Verifiable Reinforcement Learning (Zhu et al., 2019), hereafter referred to as VRL) stands out as the state-of-the-art. VRL crafts programmatic runtime shields capable of effectively rectifying all unsafe commands to consistently uphold desired safety properties (see Section 2 for more details). However, these shields exhibit limitations in the other two attributes. First, VRL’s shields involve operations with multiple high-degree polynomials, resulting in a considerable large time and memory overhead on the systems they operate within. Second, their lack of permissiveness often causes unnecessary interventions in otherwise safe commands, consequently compromising system stability. These limitations underscore opportunities for improving VRL’s shields, necessitating our work to advance the state-of-the-art.
@@
We introduce Aegis, a novel method to synthesize efficient and permissive programmatic runtime shields for neural policies. The core idea of Aegis is to formulate the seeking of a shield as a sketch-based program synthesis problem and use counterexample-guided inductive synthesis (CEGIS) (Solar-Lezama, 2008) and Bayesian optimization (BO) (Garnett, 2023) to solve it. These two techniques have been studied in program synthesis (Abate et al., 2018; Češka et al., 2019; Saad et al., 2019; Li et al., 2024) and related areas such as program repair (Ahmed et al., 2022), program translation (Zhang et al., 2021), specification inference (Hallahan et al., 2025; Klinkenberg et al., 2024) and more (Chen et al., 2024; Pailoor et al., 2024). However, Aegis is the first to combine them to enhance the efficiency and permissiveness of runtime shields, offering a novel and unique perspective to the field.
@@
Two thrusts make Aegis superior to the state-of-the-art. First, the shields synthesized by Aegis involve only linear functions or linear inequalities instead of the high-degree polynomials in VRL’s shields. This design choice reduces computational complexity and overhead, making our shields much more efficient. Second, our Bayesian optimization algorithm assigns the optimal threshold to the switching strategy by minimizing unnecessary interventions, thus making our shields more permissive than those synthesized by VRL.
@@
[…]
Novelty and Contributions
While Aegis leverages two established methodological paradigms, i.e., counterexample-guided inductive synthesis (CEGIS) and Bayesian optimization (BO), its novelty lies in how these paradigms are instantiated, customized, and tightly integrated for the specific task of synthesizing runtime shields for neural policies. Unlike prior shield synthesis works (Zhu et al., 2019; Yang et al., 2021b) or even other applications of CEGIS in broader domains (Ahmed et al., 2022; Solar-Lezama et al., 2008; Solar-Lezama, 2009; Ding et al., 2022; Edwards et al., 2024; Li et al., 2024; Pailoor et al., 2024; Zhang et al., 2021), Aegis uniquely formulates the synthesis problem as a sketch-based search over linear templates and introduces a novel refinement algorithm that leverages the maximal output admissible set to identify counterexamples and refine the shields (see Section 3.4 for more details). Furthermore, we are the first to incorporate BO into shield synthesis, tailoring it with a domain-specific objective function to optimize shields’ permissiveness by selecting appropriate switching thresholds (see Section 3.5 for more details). In addition, to the best of our knowledge, Aegis is the first to combine CEGIS and BO into a unified framework for shield synthesis, which we believe is both technically novel and offers a unique perspective on the safety assurance of neural controllers.
To summarize, our main contributions are as follows:
  • We improve both the efficiency and permissiveness of programmatic runtime shields for neural policies, addressing key limitations of the state-of-the-art.
  • We propose and implement Aegis, a novel synthesis method that formulates runtime shield generation as sketch-based program synthesis, and is the first to combine CEGIS and BO for this purpose.
  • We evaluate Aegis on eight representative control systems and show its superiority over the state-of-the-art, with significant improvements in all metrics.
@@ 2. Preliminary
Figure 2. Main control loop in a quadcopter (simplified for readability). The green and red colors highlight the neural policy and runtime shield, respectively.
@@
Modern control systems are typically implemented with digital rather than analog techniques (Lee, 2015; Zhu et al., 2019), where signals are processed discretely over fixed time intervals. Thus, the state transition of an environment can be formulated with discrete and fixed time step $\Delta t$ using Euler’s method (Hildebrand, 1987; Zhu et al., 2019). Namely, a state $s_{t}$ transitions to the subsequent state $s_{t+1}$ after time $t$ by the following equation:
@@
Here, $c_{t}$ denotes a command generated by a neural policy $\pi$, which is implemented as a neural network that takes an environment state $s_{t}$ as input, i.e., $c_{t}=\pi(s_{t})$, and $\Delta t$ is the fixed time step. Typically, neural policies are trained using Deep Reinforcement Learning (DRL) algorithms (Mousavi et al., 2018; Lillicrap et al., 2016), which optimize the policy’s parameters to maximize the expected discounted cumulative reward through interaction with the environment. Specifically, in an environment $\mathcal{E}$, a system begins at an initial state $s_{0}\in\mathcal{S}_{0}$, and its neural policy produces commands to control the system to interact with the environment until it reaches a termination state, $s_{T}$, after $T$ time steps. The interaction history is called a trajectory:
@@
The specific DRL algorithms for solving the above equation vary with different tasks and are beyond our paper’s scope. We refer interested readers to (Mousavi et al., 2018; Lillicrap et al., 2016) for more details.
@@
Figure 4. A runtime shield synthesized by Aegis for the neural policy in a quadcopter system. Line 2 (marked in red) is a linear policy that outputs safe commands, while line 3 (marked in green) identifies potentially unsafe commands.
@@ 3. Aegis
3.1. Overview
@@
Figure 6. Pendulum system (left) used as a model for the Segway personal transporter (right); image from Wikipedia (Adville, 2020). $\eta$ is the pendulum’s angle, $\omega$ is the angular velocity, and $c$ is the command from a neural policy to adjust wheel motion.
This section provides details of Aegis, with its workflow illustrated in Figure 5. Aegis begins by constructing a program sketch that consists of a linear policy and a switching strategy (Section 3.2). Then, given a complex and possibly unknown physical environment, Aegis derives its state transition function by numerical differentiation (and, 2013) (Section 3.3). Next, Aegis initializes a linear policy using the Linear-Quadratic Regulator (LQR) algorithm (Anderson and Moore, 1990) and iteratively refines it with a counterexample-guided inductive synthesis algorithm, until the linear policy is verified to be safe against the safety specifications (Section 3.4), determining the parameters of the linear policy. Following this, Aegis synthesizes the threshold of the switching strategy using a Bayesian optimization algorithm, aiming to minimize unnecessary shield interventions (Section 3.5). Finally, the synthesized components are incorporated into the program sketch to form a programmatic runtime shield. The following subsections describe each step in detail.
An Illustrative Example
To concretely demonstrate Aegis’s synthesis process, we adopt the pendulum system as a running example. This classic benchmark abstracts a rigid pendulum pivoted at the base and serves as a simplified model for many real-world control systems that require high assurance but remain difficult to verify (Zhu et al., 2019; Guo et al., 2023; Xiong and Jagannathan, 2022). In our example, we use it to model the widely-used Segway personal transporter, as shown in Figure 6. Here, the rider and the steering column are represented as a pendulum mounted on wheels. The system’s state consists of the angle $\eta$, representing the deviation of the pendulum from the vertical axis, and the angular velocity $\omega$, denoting the changing rate of this angle. A neural policy produces control commands $c$ to maintain balance by controlling the motion of the wheels, which in turn affects the pendulum’s orientation. Following Zhu et al. (Zhu et al., 2019), the system starts from a set of initial states $\mathcal{S}_{0}$:
(4) $\mathcal{S}_{0}=\left\{s\in\mathcal{S}\mid|\eta|\leq 20^{\circ}\wedge|\omega|\leq 20^{\circ}/\text{s}\right\}$
The safety property we wish to preserve is that the pendulum never falls into an unrecoverable state nor swings excessively, which can be quantitatively specified as constraining the angle $\eta$ to remain within $30^{\circ}$ and the angular velocity $\omega$ within $30^{\circ}/\text{s}$. Formally, these constraints, i.e., the set of safe states $\neg\mathcal{S}_{u}$, can be expressed as:
(5) $\neg\mathcal{S}_{u}=\left\{s\in\mathcal{S}\mid|\eta|<30^{\circ}\wedge|\omega|<30^{\circ}/\text{s}\right\}$
Figure 7. Aegis’s program sketch for the pendulum system. $\theta_{1}$, $\theta_{2}$, and $\lambda$ are the parameters to be synthesized.
As illustrated in Figure 5, the first step of Aegis involves constructing a program sketch for the runtime shield. In this example, the sketch is depicted in Figure 7. It takes as input the control command $c$ produced by the neural policy and the environment state $(\eta,\omega)$, and returns a command that satisfies safety constraints. Line 2 defines a linear policy that generates a safe command $\mathbf{K}$, where the coefficients $\theta_{1}$ and $\theta_{2}$ are to be synthesized. Line 3 specifies a switching strategy that determines when to replace the potentially unsafe command $c$ with the linear policy’s command $\mathbf{K}$, with the switching threshold $\lambda$ also to be synthesized.
Aegis then infers the unknown environment dynamics function $f(\eta,\omega,c)$, which describes how the pendulum system responds to commands from the neural policy in a given state. It approximates the dynamics as a linear time-invariant system and derives a linear approximation $f^{\prime}(\eta,\omega,c)$ using numerical differentiation (and, 2013), as shown below (more details in Section 3.3):
(6) $f^{\prime}(\eta_{t},\omega_{t},c_{t})=\begin{bmatrix}1.9027&-1\\ 1&0\end{bmatrix}\begin{bmatrix}\eta_{t}\\ \omega_{t}\end{bmatrix}+\begin{bmatrix}1\\ 0\end{bmatrix}c_{t}\approx f(\eta_{t},\omega_{t},c_{t})$
Thus, the state transition function, which describes how the pendulum system evolves from the current state $(\eta_{t},\omega_{t})$ at time $t$ to the subsequent state $(\eta_{t+1},\omega_{t+1})$ under the command $c_{t}$ from the neural policy, can be written as:
(7) $\begin{bmatrix}\eta_{t+1}\\ \omega_{t+1}\end{bmatrix}\approx\begin{bmatrix}\eta_{t}\\ \omega_{t}\end{bmatrix}+f^{\prime}(\eta_{t},\omega_{t},c_{t})\,\Delta t=\begin{bmatrix}\eta_{t}\\ \omega_{t}\end{bmatrix}+\left(\begin{bmatrix}1.9027&-1\\ 1&0\end{bmatrix}\begin{bmatrix}\eta_{t}\\ \omega_{t}\end{bmatrix}+\begin{bmatrix}1\\ 0\end{bmatrix}c_{t}\right)\Delta t$
where $\Delta t$ is the fixed time step.
Figure 8. The shield with the initialized linear policy.
With the inferred dynamics, Aegis initializes the linear policy via the Linear-Quadratic Regulator (LQR) algorithm (Anderson and Moore, 1990), which computes the policy parameters $\theta_{1}=-1.90175022$ and $\theta_{2}=0.99975035$, as shown in line 2 (highlighted in red) of Figure 8. Further computational details are provided in Section 3.4.
Then, Aegis verifies the linear policy’s safety by computing the Maximal Output Admissible Set (MOAS) (Gilbert and Tan, 1991) (see Section 3.4 for details). The MOAS identifies the initial states that can be safely controlled by the linear policy, i.e., states in $\mathcal{S}_{0}$ from which the system can avoid entering unsafe states $\mathcal{S}_{u}$. If the MOAS does not fully encompass $\mathcal{S}_{0}$, this indicates that the linear policy is unsafe, as some initial states may lead the system into unsafe regions. In such cases, the uncovered states are treated as counterexamples to guide policy refinement. For the pendulum example, in the first iteration, the boundary state with $\eta=20^{\circ}$ and $\omega=20^{\circ}/\text{s}$ lies outside the MOAS (the full MOAS expression is omitted here for brevity), demonstrating that the initial policy is insufficiently safe. Aegis then refines the policy using the algorithm described in Section 3.4, adjusting $\theta_{1}$ and $\theta_{2}$ to expand the MOAS until it fully covers $\mathcal{S}_{0}$. After several refinements, the final linear policy is obtained with $\theta_{1}=-1.91256926$ and $\theta_{2}=0.98893131$, as shown in line 2 (highlighted in red) of Figure 9.
Figure 9. The shield with the refined linear policy.
Figure 10. The final runtime shield synthesized by Aegis.
The next step is to synthesize the switching strategy, which introduces a threshold $\lambda$ that determines when to replace the command $c$ from the neural policy with the command $\mathcal{K}$ from the linear policy. If this threshold is not properly set, the system may remain unsafe, as the verified safe linear policy would not be invoked to override unsafe commands. Aegis synthesizes this threshold using a Bayesian optimization algorithm (see Section 3.5 for details), which aims to minimize safety violations while also limiting unnecessary interventions by the linear policy, thereby preserving the system’s stability. The synthesized threshold is $0.21988415$, meaning that if the commands from the neural and linear policies differ by more than $0.21988415$, the system will use the linear policy’s command instead. The final runtime shield is shown in Figure 10, where line 3 (highlighted in green) specifies the synthesized switching strategy.